# Unveiling the Sagittarius Dwarf Spheroidal Galaxy Core with Gaia DR3

### v1

@@ Abstract
The Sagittarius dwarf spheroidal galaxy provides us with the unique opportunity to study an ongoing Galactic cannibalistic event between our Milky Way Galaxy and a satellite dwarf galaxy. We present the first major membership star catalog of the Sagittarius dwarf core (400,000 sources) and Messier 54 (2000 sources) with positions, proper motions, and parallaxes from DR3, supplemented with metallicities from the Apache Point Observatory Galactic Evolution Experiment. We isolate the Sagittarius dwarf core (3.7 core radii) and Messier 54 (0.125 core radii) using literature positions. Using evolutionary sub-samples separated within a color-magnitude diagram, we analyze the structures and substructures of the Sagittarius core and infer its positional relationship with Messier 54 within 5D phase space. A sample of Milky Way stars from a similar galactic latitude was used to identify contaminants and separate member stars from the core of the Sgr dSph and Messier 54 using a Gaussian distribution in proper motion space. We present the derived proper motion, parallaxes, and metallicities for these evolutionary sub-samples while demonstrating the precision of our sample using red clump standard candles. We find a distance modulus for the Sagittarius core and Messier 54 of mag and mag, corresponding to a heliocentric distance of kpc and kpc respectively. With red clump distance analysis, our results imply there is no separation between the Sagittarius core and Messier 54.
1 Introduction
@@
The Sagittarius dwarf spheroidal galaxy (Sgr dSph, hereafter Sgr, Ibata et al., 1994, 1995), discovered nearly 30 years ago, is the closest known example of a dwarf satellite galaxy undergoing tidal disruption and accretion into a larger galaxy–the Milky Way (MW) (Johnston et al., 1995; Zhao, 1998; Johnston, 1998; Helmi et al., 1999; Bellazzini et al., 2003; Putman et al., 2004; Lewis & Ibata, 2005; Bellazzini et al., 2008; Laporte et al., 2018; Bellazzini et al., 2020; Horta et al., 2023). Dwarf galaxies are considered the building blocks of the hierarchical merging process (Gallagher & Wyse, 1994; Tosi, 2003) and, thus, are fundamental mechanisms for forming large galaxies through a process of galactic cannibalism for spiral galaxies (Hausman & Ostriker, 1978; Helmi, 2001). Tidal interactions during close encounters with large host galaxies often transform star-forming, gas-rich dwarf irregular (dIrr) galaxies with small rotating disks into classical, gas-depleted, pressure-supported dwarf spheroidals (Frinchaboy et al., 2012). In the case of Sgr, these forces from the MW’s tidal field have elongated its core toward the Galactic plane (Majewski et al., 2003), complicating observations due to its proximity of kpc below (Mucciarelli et al., 2017) and kpc behind the MW bulge (Ibata, 1999). Three pericentric passages with the MW have stripped the Sgr system, creating two massive tidal streams—a trailing arm in the south-galactic hemisphere and a leading arm in the north-galactic hemisphere–populating the MW halo with 5% of its halo M giant stars and associated dark matter particles (Majewski et al., 2003; Bellazzini et al., 2020).
@@
These pericentric passages and crossings through the MW disk have coincided with increased periods of star formation within Sgr and the MW, occurring 5.9, 1.9, and 1 Gyr ago (de Boer et al., 2015; Ruiz-Lara et al., 2020). These episodes have led to a dominance of intermediate-age stars within the Sgr core, aged around 4 to 6 Gyr, with metallicities ranging from [Fe/H] = to (Siegel et al., 2007). A younger population with [Fe/H] = formed approximately 1 Gyr ago, aligns with the most recent pericentric passage, and metal-poor old population of [Fe/H] = (Siegel et al., 2007; Alfaro-Cuello et al., 2019). Although material from Sgr is dispersed across a wide range of distances (10-100 kpc) due to the tidal disruption, parts of the main body remain embedded within these tidal streams, carrying stars and dark matter lost during this process (Hasselquist et al., 2017; Fernández-Trincado et al., 2021). While understanding the tidal streams of Sgr are crucial for probing the MW gravitational potentials and changes within the disk, understanding the core dynamics, physical properties, and kinematics, of the Sgr main body is essential for deciphering the evolution and nature of the Sgr system. Investigating these properties through galactic archaeology—studying galactic structure and evolution by examining the ages, chemical compositions, and distances of stellar populations—offers insight into Sgr and its relationship with M54.
@@
Following the release of DR3, Gaia Collaboration et al. (2021) probed the structure of the LMC and SMC, showcasing improvements from Gaia Collaboration et al. (2018) where they previously determined the kinematics, including mean proper motion (, ) and parallax of 75 Galactic globular clusters, the Large and Small Magellanic clouds (LMC and SMC respectively), and nine dwarf spheroidal galaxies including the Sgr core within DR2. They found that photometry formerly affected by background issues caused by high stellar density in central areas of dwarfs were reduced, there was a two-fold reduction in proper motion uncertainty, and an overall increase in the precision of astrometry and photometry with significantly reduced systematic effects from DR2. APOGEE DR17 (Abdurro’uf et al., 2022a) contains measurements for stars from dwarf satellite galaxies, open clusters, and globular clusters (GCs) identified from procedures detailed in Zasowski et al. (2013), including Sgr and M54. The mixture of these two sky surveys allows us to constrain Sgr and M54 member stars from the MW disk to analyze the dynamics within 5D phase space.
In this work, we describe membership selection of the Sgr core and M54 using DR3 and APOGEE DR17, inspired by the methodology within Gaia Collaboration et al. (2018, 2021). We describe the data selection procedure in Section 2. In Section 3 we describe the spatial separation process, removing background MW contaminants, and isolating evolutionary areas of the Sgr core and M54. In Section 4, we use our sample of red clump (RC) stars to estimate a distance to Sgr and M54 and present our result from our distance determination. In Section 5, we present results from 5D phase space analysis of the Sgr core and M54, compare our distance determinations from the RC with prior studies. We compare our membership sample to previous work and discuss the efficacy of our sample for future studies. In Section 6, we summarize our conclusions.
2 Data
2.1 Gaia DR3
@@
We extract a sample from the DR3 catalog based on specific spatial criteria, where , are the accepted center of the Sgr core from literature in right ascension and declination respectively using the coordinates () = (283.75∘, -30.46∘) (Majewski et al., 2013), with a radius of . This initial large selection radius is greater than defined in Majewski et al. (2013) as the core of Sgr has a defined ellipticity of (Majewski et al., 2013). We constrain our selection to sources that have at least a five-parameter astrometric solution. This selection can be reproduced in full using the ADQL query within Appendix 8.1. The resulting sample contains 4,378,343 objects.
2.2 SDSS/APOGEE DR17
@@
The APOGEE spectra were processed using the APOGEE reduction pipeline (Nidever et al., 2015), and key parameters were derived through the APOGEE Stellar Parameters and Chemical Abundances Pipeline (ASPCAP; Garc´ıa Pérez et al., 2016). Further details on the spectral analysis process can be found in Shetrone et al. (2015); Smith et al. (2021). The final data release for APOGEE-2 (DR17; Abdurro’uf et al., 2022b) includes approximately 734,000 stars, encompassing all observations from both spectrographs collected between August 2011 and January 2021. A comprehensive overview of the quality and parameter limitations of APOGEE DR17 is available in Abdurro’uf et al. (2022b).
3 Methodology
3.1 Spatial Selection
From our DR3 query, we compute the effective radius of the Sgr core from flattened coordinates defined in Section 8.2, using , the accepted center of the Sgr core from literature (in right-ascension and declination), the ellipticity, and the position angle, as defined from Majewski et al. (2013). This spatial selection includes the globular cluster Messier 54 (M54; NGC6715), which has a tidal radius of 7.5 arcmin (0.125∘) (Trager et al., 1995).
@@ 3.1 Spatial Selection
To refine our samples and reduce contamination from foreground Milky Way stars, we applied a modified version of the methods described by Gaia Collaboration et al. (2018) and Gaia Collaboration et al. (2021).
@@
This contamination is evident in Gaia Collaboration et al. (2018, figure 3) within the color-magnitude diagrams (CMDs) at . We reproduce the figure using our sample defined here, where background objects are also apparent in the top row of CMDs for the Sgr core (left) and M54 (right) within Figure 3.
3.2 Evolutionary Features
@@
Before cleaning of Milky Way contaminants, we isolate evolutionary features in our CMDs for Sgr dSph core and M4. Based on Figure 3 in Gaia Collaboration et al. (2018), we loosely divided our CMDs into three color regions defined as dex, dex dex, and mag (blue, green, and red regions). The blue region appears nearly free of Milky Way contamination, whereas the green and red zones appear heavily contaminated, yet the features of the Sgr and M54 stand out, particularly when viewing stars that have measured [Fe/H] values from APOGEE DR17 in the top panel of Figure 3. We identify the following distinct evolutionary phases within the Sgr core based on Figure 2 from Gaia Collaboration et al. (2021): a young population extending beyond the main sequence turn-off (Young); the blue main sequence (BL MS); the red giant branch (RGB); the asymptotic giant branch (AGB), including long-period variables; the RR Lyrae (RR Lyr) region; the blue horizontal branch (BHB), which also includes RR Lyrae stars; the blue loop (BL), encompassing classical Cepheids; and the red clump (RC). Furthermore, since Sgr is an evolved, disrupted dwarf spheroidal, unlike the Large and Small Magellanic clouds in Gaia Collaboration et al. (2021), we additionally reference the methodology for selecting evolutionary features in de Boer et al. (2012, figure 4) due to the similarity between the Sgr dSph system and the Sculptor dwarf spheroidal system.
@@
Our defined areas are shown in Figure 3, with coordinates for the polygonal areas in Sgr core CMD shown in Table 1 and for M54 within Table 2. There are unassigned regions with the CMD diagrams, as these areas are prevalent of mixing of evolutionary areas, heavy Milky Way contamination, and/or are affected by blended stars. The most notable area of blended stars is within the green and red zones where we predict the main sequence, main sequence turn off, and sub-giant branch would lie, but cannot be disentangled. We note that the CMD for M54 excludes the region for young stars and the blue loop, as these features were not visible. The RGB and AGB within the Sgr core and M54 are noticeable features when viewing the CMDs colored by available [Fe/H] values in Figure 3, where both objects feature distinct RGB and AGB branches of differing metallicities (agreeing with the two branches identified in Minelli et al., 2023, one more metal-poor and one more metal-rich).
| Evolutionary Feature | Polygon Coordinates |
| Young | [0.7, 17.8], [-0.5, 17.8], [0.2, 15], [0.7, 15], [0.7, 17.8] |
| [2.6, 15], [2.0, 16.2], [1.6, 17.8], [3.0, 19.5] | |
| AGB | [1.2, 15], [1.2, 13], [1.5, 13], [4.0, 13], [4.0, 15], [1.2, 15] |
| BL | [1, 16], [0.7, 15], [0.5, 12], [1.2, 12], [1.2, 13], [1.2, 15] |
| RC | [1.1, 18.1], [1.1, 17.8], [1.4, 17.8], [1.4, 18.1], [1.1, 18.1] |
| Evolutionary Feature | Polygon Coordinates |
| MS | [-0.5, 20.5], [-0.5, 20], [2.5, 20], [2.5, 20.5], [-0.5, 20.5] |
| [2.6, 15], [2, 16.2], [1.6, 17.8], [3, 19.5] | |
| AGB | [1.2, 15], [1.2, 14], [1.5, 14], [4.4, 14], [4.4, 15], [1.2, 15] |
| RC | [1.1, 18.1], [1.1, 17.8], [1.5, 17.8], [1.5, 18.1], [1.1, 18.1] |
@@
3.3 Minimizing Milky Way Contamination
The sample of objects obtained in the previous section of the Sgr core and M54 are still heavily contaminated by foreground and background contaminants. To clean our samples of these foreground and background contaminants, we split our larger sample into the identified evolutionary features from our contaminated CMDs in Section 3.2, along with the non-identified evolutionary areas.
@@ 3.3 Minimizing Milky Way Contamination
In each corresponding histogram of and for each defined evolutionary feature and within our identified contaminated areas, there exist two distinct peaks, indicating two distinct populations within our sample as seen in the top and right panel in Figure 2. Using reported literature values from Gaia Collaboration et al. (2018) for the Sgr core within Table C.2, we identify one of these peaks as being Sgr member stars with a and .
@@
To confirm that the other peak present are Milky Way contaminants, we pulled another sample from DR3 at the same galactic latitude of the Sgr core, within a radius, but with an offset of of Galactic longitude, where (referencing and for the Sgr core; Majewski et al. (2013).) We employ the same quality cuts as in Section 3.1, and calculate the respective mean proper motions in the right ascension and declination direction finding, and . These calculated values align with the tallest peak within our histograms as featured in Figure 2.
To remove these Milky Way contaminants, we employ a Gaussian distribution fit taking the peak at the literature values of the proper motion in the right ascension () and declination () direction from Gaia Collaboration et al. (2018), with a , on areas of identifiable evolutionary features within the Sgr CMD which display the tallest spike of Sgr members.
We then apply a Gaussian fit to stars not identified within our evolutionary features to provide a general cleaning to limit as many Milky Way contaminants within those regions as well, but increasing the two-fold to for a more liberal cleaning. We remove a total of 1,467,943 likely background Milky Way contaminants from the Sgr core, leaving a sample of 515,689 member stars of the Sgr core. Figure 2 visualizes the separation of Sgr-identified members (in red) in the left-most panel and identified MW contamination (in blue).
For M54, there are fewer MW background contaminants visible (in part because our sample is smaller), yet we apply the same cleaning method as before for both the evolutionary subsamples and non-identified evolutionary subsamples. We reuse our prior literature values from Gaia Collaboration et al. (2018) along with the same as there are no prior literature measurements for the proper motion for M54. We visualize the identified Gaussian fits to our histogram evolutionary subsamples in the right-most subplot in Figure 2. We remove a total of 3,340 likely background Milky Way contaminants from M54, leaving a sample of 4,512 member stars of M54.
@@
With our final samples for the Sgr core and M54, we present an overview of the mean astrometric parameters and number of stars within each evolutionary subsample in Table 3. The mean parallax of the Sgr core from Monaco et al. (2004); Minelli et al. (2023) (calculated from the heliocentric distance) was mas. Most of our parallax measurements lie somewhat near this value of mas, however reliable distances cannot be obtained by inverting the parallax beyond kpc, causing the discrepancy in our calculated distance measurements.
4 Red Clump Distance Determination

### v2

@@ Abstract
The Sagittarius dwarf spheroidal galaxy provides us with the unique opportunity to study an ongoing Galactic cannibalistic event between our Milky Way Galaxy and a satellite dwarf galaxy. Understanding this event crucially requires memberships and high-precision metallicities. Here, we present the first major membership star catalog of the Sagittarius dwarf core (140,000 sources) and Messier 54 (2000 sources) with positions, proper motions, and parallaxes from DR3, supplemented with metallicities from the Apache Point Observatory Galactic Evolution Experiment. We initially isolate the Sagittarius dwarf core and Messier 54 spatially from prior literature positions. Using evolutionary sub-samples separated within a color-magnitude diagram, we analyze the substructures of the Sagittarius core and infer its positional relationship with Messier 54 within 5D phase space. A sample of Milky Way stars from a similar galactic latitude were used to identify contaminants and separate member stars from the core of the Sgr dSph and Messier 54 using a Gaussian Mixture Model. We present the derived proper motion, parallaxes, and metallicities for these evolutionary sub-samples while demonstrating the precision of our sample using red clump standard candles. We find a distance modulus for the Sagittarius core and Messier 54 of mag and mag, corresponding to a heliocentric distance of kpc and kpc respectively. With red clump distance analysis, our results imply there is no separation between the Sagittarius core and Messier 54. Finally, we describe the metallicity distributions of the evolved stars within these two systems, finding evidence for the infall scenario.
I Introduction
@@
The Sagittarius dwarf spheroidal galaxy (Sgr dSph, hereafter Sgr, Ibata et al., 1994, 1995), discovered nearly 30 years ago, is the closest known example of a dwarf satellite galaxy undergoing tidal disruption and accretion into a larger galaxy–the Milky Way (MW) (Johnston et al., 1995; Zhao, 1998; Johnston, 1998; Helmi et al., 1999; Bellazzini et al., 2003; Putman et al., 2004; Lewis & Ibata, 2005; Bellazzini et al., 2008; Laporte et al., 2018; Bellazzini et al., 2020; Horta et al., 2023). Dwarf galaxies are considered the building blocks of the hierarchical merging process (Gallagher & Wyse, 1994; Tosi, 2003; Tolstoy et al., 2009; Helmi, 2020) and, thus, are fundamental mechanisms for forming large galaxies through a process of galactic cannibalism for spiral galaxies (Hausman & Ostriker, 1978; Helmi, 2001). Tidal interactions during close encounters with large host galaxies often transform star-forming, gas-rich dwarf irregular (dIrr) galaxies with small rotating disks into classical, gas-depleted, pressure-supported dwarf spheroidals (Frinchaboy et al., 2012). In the case of Sgr, these forces from the MW’s tidal field have elongated its core toward the Galactic plane (Majewski et al., 2003), complicating observations due to its proximity of kpc below (Mucciarelli et al., 2017) and kpc behind the MW bulge (Ibata, 1999). Three pericentric passages with the MW have stripped the Sgr system, creating two massive tidal streams—a trailing arm in the south-galactic hemisphere and a leading arm in the north-galactic hemisphere–populating the MW halo with 5% of its halo M giant stars and associated dark matter particles (Majewski et al., 2003; Bellazzini et al., 2020).
@@
These pericentric passages and crossings through the MW disk have coincided with increased periods of star formation within Sgr and the MW, occurring 5.9, 1.9, and 1 Gyr ago (de Boer et al., 2015; Ruiz-Lara et al., 2020). These episodes have led to a dominance of intermediate-age stars within the Sgr core, aged around 4 to 6 Gyr, with metallicities ranging from [Fe/H] = to (Siegel et al., 2007). A younger, metal-rich population was first detected by Bonifacio et al. (2000) and the complex star formation history of Sgr dSph was further studied by Bonifacio et al. (2004); Sbordone et al. (2007). This near-solar metal-rich population formed approximately 1 Gyr ago, aligns with the most recent pericentric passage, and the metal-poor old population of [Fe/H] = corresponds to pre-pericentric passages (Bonifacio et al., 2004; Alfaro-Cuello et al., 2019). Although material from Sgr is dispersed across a wide range of distances (10-100 kpc) due to the tidal disruption, parts of the main body remain embedded within these tidal streams, carrying stars and dark matter lost during this process (Hasselquist et al., 2017; Fernández-Trincado et al., 2021). While understanding the tidal streams of Sgr are crucial for probing the MW gravitational potentials and changes within the disk, understanding the core dynamics, physical properties, and kinematics, of the Sgr main body is essential for deciphering the evolution and nature of the Sgr system. Investigating these properties through galactic archaeology—studying galactic structure and evolution by examining the ages, chemical compositions, and distances of stellar populations—offers insight into Sgr and its relationship with M54.
@@
Following the release of DR3, Gaia Collaboration et al. (2021) probed the structure of the LMC and SMC, showcasing improvements from Gaia Collaboration et al. (2018a) where they previously determined the kinematics, including mean proper motion (, ) and parallax of 75 Galactic globular clusters, the Large and Small Magellanic clouds (LMC and SMC respectively), and nine dwarf spheroidal galaxies including the Sgr core within DR2. They found that photometry formerly affected by background issues caused by high stellar density in central areas of dwarfs were reduced, there was a two-fold reduction in proper motion uncertainty, and an overall increase in the precision of astrometry and photometry with significantly reduced systematic effects from DR2. APOGEE DR17 (Abdurro’uf et al., 2022a) contains measurements for stars from dwarf satellite galaxies, open clusters, and globular clusters (GCs) identified from procedures detailed in Zasowski et al. (2013), including Sgr and M54. The mixture of these two sky surveys allows us to constrain Sgr and M54 member stars from the MW disk to analyze the dynamics within 5D phase space.
In this work, we describe membership selection of the Sgr core and M54 using DR3 and APOGEE DR17, inspired by the methodology within Gaia Collaboration et al. (2018a, 2021). We describe the data selection procedure in Section II. In Section III we describe the spatial separation process, removing background MW contaminants, and isolating evolutionary areas of the Sgr core and M54. In Section IV, we use our sample of red clump (RC) stars to estimate a distance to Sgr and M54 and present our result from our distance determination. In Section V, we present results from 5D phase space analysis of the Sgr core and M54, compare our distance determinations from the RC with prior studies. We compare our membership sample to previous work and discuss the efficacy of our sample for future studies. In Section VI, we summarize our conclusions.
II Data
II.1 Gaia DR3
@@
We extract a sample from the DR3 catalog based on specific spatial criteria, where , are the accepted center of the Sgr core from literature in right ascension and declination respectively using the coordinates () = (283.75∘, -30.46∘) (Majewski et al., 2003), with a radius of . This initial large selection radius is greater than defined in Majewski et al. (2003) as the core of Sgr has a defined ellipticity of (Majewski et al., 2003). We constrain our selection to sources that have at least a five-parameter astrometric solution. This selection can be reproduced in full using the ADQL query within Appendix A. The resulting sample contains 4,378,343 objects.
II.2 SDSS/APOGEE DR17
@@
The APOGEE spectra were processed using the APOGEE reduction pipeline (Nidever et al., 2015), and key parameters were derived through the APOGEE Stellar Parameters and Chemical Abundances Pipeline (ASPCAP; García Pérez et al., 2016). Further details on the spectral analysis process can be found in Shetrone et al. (2015); Smith et al. (2021). The final data release for APOGEE-2 (DR17; Abdurro’uf et al., 2022b) includes approximately 734,000 stars, encompassing all observations from both spectrographs collected between August 2011 and January 2021. A comprehensive overview of the quality and parameter limitations of APOGEE DR17 is available in Abdurro’uf et al. (2022b).
III Methodology
III.1 Spatial Selection
From our DR3 query, we compute the effective radius of the Sgr core from flattened coordinates defined in Section B, using , the accepted center of the Sgr core from literature (in right-ascension and declination), the ellipticity, and the position angle, as defined from Majewski et al. (2013). This spatial selection includes the globular cluster Messier 54 (M54; NGC6715), which has a tidal radius of 7.5 arcmin (0.125∘) (Trager et al., 1995).
@@ III.1 Spatial Selection
To refine our samples and reduce contamination from foreground Milky Way stars, we applied a modified version of the methods described by Gaia Collaboration et al. (2018a) and Gaia Collaboration et al. (2021).
@@
This contamination can be seen in Gaia Collaboration et al. (2018a, Figure 3) and Gaia Collaboration et al. (2018b, Figure 1) at around , increasing as we move into the redder color zone. We can see this within our sample within Figure 7.
@@
We reproduce the figure using our sample defined here, where background objects are also apparent in the top row of CMDs for the Sgr core (left) and M54 (right), within Figure 3.
III.2 Evolutionary Features
@@
Before cleaning of Milky Way contaminants, we isolate evolutionary features in our CMDs for Sgr dSph core and M54. Based on Figure 3 in Gaia Collaboration et al. (2018a), we loosely divided our CMDs into three color regions defined as dex, dex dex, and mag (blue, green, and red regions; see Figure 7). The blue region appears nearly free of Milky Way contamination, whereas the green and red zones appear heavily contaminated, yet the features of the Sgr and M54 stand out, particularly when viewing stars that have measured [Fe/H] values from APOGEE DR17 in the top panel of Figure 3. We identify the following distinct evolutionary phases within the Sgr core based on Figure 2 from Gaia Collaboration et al. (2021): a young population extending beyond the main sequence turn-off (Young); the blue main sequence (BL MS); the red giant branch (RGB); the asymptotic giant branch (AGB), including long-period variables; the RR Lyrae (RR Lyr) region; the blue horizontal branch (BHB), which also includes RR Lyrae stars; the blue loop (BL), encompassing classical Cepheids; and the red clump (RC). Furthermore, since Sgr is an evolved, disrupted dwarf spheroidal, unlike the Large and Small Magellanic clouds in Gaia Collaboration et al. (2021), we additionally reference the evolutionary subsamples in de Boer et al. (2012, Figure 3) as both Fornax and Sgr are dwarf spheroidals.
@@
Our defined areas are shown in Figure 3, with coordinates for the polygonal areas in Sgr core CMD shown in Table 3 and for M54 within Table 4. There are unassigned regions with the CMD diagrams, as these areas are prevalent of mixing of evolutionary areas, heavy Milky Way contamination, and/or are affected by blended stars. The most notable area of blended stars is within the green and red zones where we predict the main sequence, main sequence turn off, and sub-giant branch would lie, but cannot be disentangled. We note that the CMD for M54 excludes the region for young stars and the blue loop, as these features were not visible. The RGB and AGB within the Sgr core and M54 are noticeable features when viewing the CMDs colored by available [Fe/H] values in Figure 3, where both objects feature distinct RGB and AGB branches of differing metallicities (see Boecker et al. (2020) for M54 and Minelli et al. (2023) for Sgr dSph core).
@@
III.3 Minimizing Milky Way Contamination
The sample of objects obtained in the previous section of the Sgr core and M54 are still heavily contaminated by foreground and background contaminants. To clean our samples of these foreground and background contaminants, we split our larger sample into the identified evolutionary features from our contaminated CMDs in Section III.2, along with the non-identified evolutionary areas.
@@ III.3 Minimizing Milky Way Contamination
In each corresponding histogram of and for each defined evolutionary feature and within our identified contaminated areas, there exist two distinct peaks, indicating two distinct populations within our sample as seen in Figure 2. Using reported literature values from Gaia Collaboration et al. (2018a) for the Sgr core within Table C.2, we identify one of these peaks as being Sgr member stars with a and .
@@
To confirm that the other peak present are Milky Way contaminants, we pulled two samples from DR3 at the same galactic latitude of the Sgr core, within a radius, but with an offset of of Galactic longitude, where (referencing and for the Sgr core; Majewski et al. (2003).) We employ the same quality cuts as in Section III.1, and calculate the respective mean proper motions in the right ascension and declination direction finding, and . These calculated values align with the tallest peak within our histograms as featured in Figure 2.
To remove these Milky Way contaminants, we employ a Gaussian Mixture Model (GMM) implemented using the Scikit-learn package (Pedregosa et al., 2011) to disentangle our member stars from each system from the MW contaminants. A GMM is a unsupervised machine learning clustering technique where the probabilistic model assumes that the data is generated from a mixture of two or more Gaussian distributions, each with its own weight, mean and variance Ivezić et al. (2020). By providing an initial guess of the number of components, an Expectation-Maximization (EM) Algorithm can be used to find various parameters of the Gaussian components and is repeated till convergence. This also provides a probability value for each data point, dictating which Gaussian component it belongs to.
Utilizing the literature values of the proper motion from literature values of the proper motion in the right ascension () and declination () direction from Gaia Collaboration et al. (2018a), we apply a GMM using these values as our initial guesses for the peak of our member stars on the areas of identifiable evolutionary features within the Sgr CMD, along with defining 2 number of components. We accept any stars that report a 95% or higher probability of belonging to the Gaussian identified belonging to Sgr. We report the FWHM values for the high probability members of the Sgr core as and . We perform a different process to areas outside identified evolutionary features, since the GMM cannot disentangle Sgr’s and from the MW’s and to a high probability since the MW is much more prevalent in these areas. Instead, we translate the FWHM measurements from our GMM model to an ellipse (rotated to account for Sgr’s intrinsic dispersion) and spatially cut out stars that match Sgr’s and distribution. If we were to adopt a GMM, our sample would be heavily skewed to selecting MW contamination stars, not Sgr stars. Figure 2 visualizes the separation of Sgr-identified members (in pink) in the left-most panel and identified MW contamination (in purple). Our final sample includes 144596 identified Sgr member stars, 95199 of which are high-probability Sgr members through the GMM, which are shown in Table 2.
For M54, there are fewer MW background contaminants visible (due to the high density of M54Sgr stars), yet we apply the same cleaning method as before for both the evolutionary subsamples and unidentified evolutionary subsamples. We use literature values from An et al. (2024) ( and ) corroborated with the calculated and of M54 members from APOGEE DR17. These values provide an starting point for our GMM, and we again define an initial guess of 2 Gaussian components. We report the FWHM values for the high probability members of M54 as FWHM and FWHM. We visualize the high probability identified members of M54 in the right subplot in Figure 2. Our final sample includes 2638 identified member stars of M54, which are shown in Table 2.
| Sgr dSph Sample | ||||||
| DR3 ID | RA | Dec | GMM Probability | |||
| deg | deg | mas | ||||
| 6736241413019624704 | 282.828419 | -31.592158 | -2.547946 | -1.413759 | 0.122520 | 0.988925 |
| 6736241447378768128 | 282.840249 | -31.589329 | -2.829347 | -1.473527 | -0.084372 | 0.988629 |
| 6736253060970356864 | 282.741142 | -31.566766 | -2.740621 | -1.473031 | 0.226700 | 0.990030 |
| M54 Sample | ||||||
| DR3 ID | RA | Dec | GMM Probability | |||
| 6760412458169160320 | 283.811137 | -30.599297 | -2.431466 | -1.168130 | -0.123008 | 0.976058 |
| 6760412462477482112 | 283.812538 | -30.600012 | -2.883681 | -1.446605 | -1.756710 | 0.979323 |
| 6760412492527580032 | 283.824930 | -30.587363 | -2.463567 | -1.460903 | -0.179012 | 0.975957 |
Note. Random sample of member stars from the Sgr core and M54. Stars with a GMM Probability indicated with NaN (not a number) were selected spatially using the ellipse method. The full version of these tables are available electronically in full online.
@@
With our final samples for the high probability members of the Sgr core and M54, we present an overview of the mean astrometric parameters and number of stars within each evolutionary subsample in Table 1. The mean parallax of the Sgr core from Monaco et al. (2004); Minelli et al. (2023) (calculated from the heliocentric distance) was mas. Most of our parallax measurements lie somewhat near this value of mas, however reliable distances cannot be obtained by inverting the parallax beyond kpc, causing the discrepancy in our calculated distance measurements.
IV Red Clump Distance Determination